OpenStudy (anonymous):

Find all theta in the interval [0, 2pi) that satisfy the equation. 2cos(theta)tan(theta)+tan(theta)=0

OpenStudy (mandre):

2cos(theta)tan(theta) = -tan(theta) 2cos(theta) = -1 (divide both sides by tan(theta)) cos(theta)= -1/2 \[\theta=\cos ^{-1}(-\frac{ 1 }{ 2 })\]

OpenStudy (ybarrap):